# ClinNoteAgents: An LLM Multi-Agent System for Predicting and Interpreting Heart Failure 30-Day Readmission from Clinical Notes

### v1

@@ CONCLUSION
The study introduced ClinNoteAgents, an LLM–based multi-agent framework that operationalizes two core tasks: (1) HF readmission risk-factor mining and (2) HF readmission risk prediction. By transforming unstructured discharge notes into structured representations of clinical measurements, social determinants of health, and diagnoses, the system provides an end-to-end pipeline that directly supports predictive modeling and risk-factor analysis. For the readmission-factor mining task, the coordinated extraction and normalization agents recovered clinically relevant information with high fidelity, standardized heterogeneous medical term expressions, and enabled downstream statistical association analysis that would otherwise have been infeasible on raw free-text. These results demonstrate the feasibility of using LLMs to generate analyzable structural representations in limited structured EHR environments. For the readmission prediction task, abstractions produced by the LLM summarizer preserved most predictive signals from raw discharge notes despite substantial text compression. This indicates that the summarizer can distill long and noisy clinical narratives into concise representations that reduce computational cost while retaining clinical utility.
Taken together, ClinNoteAgents provides a unified framework that enables scalable extraction, normalization, and summarization of free-text documentation to support HF readmission risk modeling. The system produces structured, interpretable abstractions compatible with rule-based pipelines, temporal reasoning modules, and clinical decision-support tools. Future work will extend ClinNoteAgents to multi-institutional datasets, integrate domain-specialized clinical LLMs, and incorporate causal-inference modules to characterize diagnosis-specific pathways of readmission risk. The framework ultimately aims to support more reliable, interpretable, and equitable clinical decision-making across diverse healthcare settings.

### v2

@@ CONCLUSION
The study introduced ClinNoteAgents, an LLM–based multi-agent framework that operationalizes two core tasks: (1) HF readmission risk-factor mining and (2) HF readmission risk prediction. By transforming unstructured discharge notes into structured representations of clinical measurements, social determinants of health, and diagnoses, the system provides an end-to-end pipeline that directly supports predictive modeling and risk-factor analysis. For the readmission-factor mining task, the coordinated extraction and normalization agents recovered clinically relevant information with high fidelity, standardized heterogeneous medical term expressions, and enabled downstream statistical association analysis that would otherwise have been infeasible on raw free-text. These results demonstrate the feasibility of using LLMs to generate analyzable structural representations in limited structured EHR environments. For the readmission prediction task, abstractions produced by the LLM summarizer preserved most predictive signals from raw discharge notes despite substantial text compression. This indicates that the summarizer can distill long and noisy clinical narratives into concise representations that reduce computational cost while retaining clinical utility. Taken together, ClinNoteAgents provides a unified framework that enables scalable extraction, normalization, and summarization of free-text documentation to support HF readmission risk modeling. The system produces structured, interpretable abstractions compatible with rule-based pipelines, temporal reasoning modules, and clinical decision-support tools. In a clinical decision-support setting, ClinNoteAgents could be integrated into the EHR to generate readmission risk predictions alongside extracted risk factors and structured note summaries. Outputs would remain solely for advisory purposes and require clinician verification to mitigate hallucination and omission risks.
ACKNOWLEDGMENT
This research was partially supported by the internal funds and GPU servers provided by School of Nursing’s Center for Data Science of Emory University, Computer Science Department of Emory University, the US National Science Foundation under Award Numbers 2442172, 2312502, 2319449, and the US National Institute of Health under Award Numbers K25DK135913, RF1NS139325, R01DK143456 and U18DP006922.